# Diffusion Attribution Score: Evaluating Training Data Influence in Diffusion Model

### v2

@@ 4.3 Evaluation for Speed Up techniques
In this section, we detail the experiments conducted to evaluate the effectiveness of the speed-up techniques applied in computing the DAS. The diffusion models used in our experiments vary significantly in complexity, with parameter counts of 35.7M, 118.8M, and 25.5M respectively. These large dimensions pose considerable challenges in calculating the attribution score efficiently. To address this, we implemented speed-up techniques as discussed in Section 3.3.
Normalization. The first technique we evaluated is the normalization of gradients and residuals, as proposed in Eq. 17. Given that the data attribution method in Eq. 16 targets model output at a specific timestep, we advocated for the normalization of gradients and residuals across all generation processes before averaging. This approach is intended to stabilize the variability in gradient magnitudes across timesteps, thereby enhancing computational efficiency and accuracy. The specifics of this normalization process are detailed in Eq. 17.
Due to space constraints, the results of these evaluations in this section are reported in the Appendix F.4. As shown in Table 4, the implementation of normalization improved the performance of both DAS and D-TRAK.
Number of timesteps. When computing the DAS as defined in Eq. 18, obtaining gradients at more timesteps can potentially improve performance due to more comprehensive averaging, but it also necessitates more back-propagation steps. Consequently, there is a notable trade-off between effectiveness and computational efficiency. To explore this, we considered the extreme case of computing gradients for all 1000 timesteps—a scenario that is computationally intensive, especially for large projection dimensions and extensive datasets. We conducted an experiment using the CIFAR-2 dataset with a projection dimension $k=4096$. The results, presented in Table 5, demonstrate that increasing the number of timesteps correlates with improved LDS outcomes. Notably, our findings also indicate that setting the number of timesteps to 100, or even 10, can achieve performance comparable to that of 1000 timesteps, but with significantly reduced computational demands. Based on these insights, in subsequent experiments, we will default to considering 10 and 100 timesteps to balance performance with computational efficiency.
Projection. We applied the projection technique discussed in Section 3.3 to reduce the dimension of gradients. We conducted experiments to explore the relationship between the projection dimensions $k$ and LDS performance. As suggested by Johnson & Lindenstrauss (1984), increasing the projection dimension tends to preserve inner products more effectively, though at a higher computational cost. Experiments in Figure 2 illustrate that as the dimension increases, the LDS scores for both D-TRAK and DAS initially rise and then gradually plateau, as anticipated. Results depicted in Figure 2 show that as the projected dimension increases, the LDS scores for both D-TRAK and DAS initially rise and then gradually plateau. Based on these findings, we have established $k=32768$ as the default setting for subsequent experiments.
Compress Model Parameters. We also explored techniques to reduce the gradient dimension at the model level. Specifically, we conducted experiments focusing solely on using the up-block of the U-Net to compute gradients. The results, as shown in Table 6, indicate that using only the up-block can achieve competitive performance compared to the full model configuration. Additionally, experiments on ArtBench-2 and ArtBench-5, which involved fine-tuning a Stable Diffusion model with LoRA, further demonstrated the effectiveness of this approach.
@@
Candidate Training Sample. Another technique to speed up the process involves reducing the number of training samples considered. We conducted an experiment using CLIP to select the training samples most similar to the target sample. Specifically, we generated a candidate dataset comprising 1,000 training samples that CLIP identified as most similar to the generated or validated image on CIFAR-2. We then computed the attribution scores for this candidate set, assigning a score of 00 to all other samples, and calculated the LDS. The results, detailed in Table 7, validate the efficacy of this method.
Table 3: LDS (%) on CelebA with timesteps (10 or 100)
Results on CelebA
Method Validation Generation
10 100 10 100
Raw pixel (dot prod.) 5.58$\pm$0.73 -4.94$\pm$1.58
Raw pixel (cosine) 6.16$\pm$0.75 -4.38$\pm$1.63
CLIP similarity (dot prod.) 8.87$\pm$1.14 2.51$\pm$1.13
CLIP similarity (cosine) 10.92$\pm$0.87 3.03$\pm$1.13
TracInCP 5.14$\pm$0.75 4.89$\pm$0.86 5.18$\pm$1.05 4.50$\pm$0.93
GAS 5.44$\pm$0.68 5.19$\pm$0.64 4.69$\pm$0.97 3.98$\pm$0.97
Journey TRAK / / 6.53$\pm$1.06 10.87$\pm$0.84
Relative IF 11.10$\pm$0.51 19.89$\pm$0.50 6.80$\pm$0.77 14.66$\pm$0.70
Renorm. IF 11.01$\pm$0.50 18.67$\pm$0.51 6.74$\pm$0.82 13.24$\pm$0.71
TRAK 11.28$\pm$0.47 20.02$\pm$0.47 7.02$\pm$0.89 14.71$\pm$0.70
D-TRAK 22.83$\pm$0.51 28.69$\pm$0.44 16.84$\pm$0.54 21.47$\pm$0.48
DAS 29.38$\pm$0.51 33.79$\pm$0.23 28.73$\pm$0.49 30.68$\pm$0.31
4.4 Main Experiment
@@
In this section, we evaluate the performance of the Diffusion Attribution Score (DAS) against existing attribution baselines that can be applied in our experimental settings, following methodologies similar to those described by Zheng et al. (2024). We limit the use of techniques to projection only, omitting others like normalization to ensure fair comparisons across different attribution methods. Our primary focus is on post-hoc data attribution methods, which are applied after the model’s training is complete. These methods are categorized into similarity-based, gradient-based (without kernel), and gradient-based (with kernel) approaches, with detailed explanations provided in Appendix F.3. The outcomes are documented in Tables 1, 2, and 3, where DAS consistently outperforms existing methods across all datasets.
@@
Compared to D-TRAK, DAS shows substantial improvements. For instance, on a validation set utilizing 100 timesteps, DAS achieves improvements of +9.33% on CIFAR-2, +8.39% on ArtBench-2, and +5.1% on CelebA. In the generation set, the gains continue with +5.01% on CIFAR-2, +5.78% on ArtBench-2, and +9.21% on CelebA. Notably, DAS also achieves significant improvements on larger datasets like CIFAR10 and ArtBench5, outperforming D-TRAK by +12.07% and +10.21% on their respective validation sets.
@@
Analysis of results. Other methods generally underperform on larger datasets such as ArtBench5 and CIFAR10 compared to smaller datasets like CIFAR2 and ArtBench2. Conversely, our method performs better on ArtBench5 than on ArtBench2. Remarkably, our findings suggest that while with more timesteps for calculating gradients generally leads to a better approximation of the expectation $\mathbb{E}_{t}$, DAS, employing only a 10-timestep computation budget, still outperforms D-TRAK, which uses a 100-timestep budget in most cases, which underscores the effectiveness of our approach. Additionally, the modest improvements on CIFAR10 and CelebA may be attributed to the LDS setup for these datasets, which employs only one random seed per subset for training a model, whereas other datasets utilize three random seeds, potentially leading to inaccuracies in LDS evaluation.
Another observation is that DAS performs better on the validation set than on the generation set. This could indicate that the quality of generated images may have a significant impact on data attribution performance. However, further investigation is needed to validate this hypothesis.
To more intuitively assess the faithfulness of DAS, we also conduct an toy example, that generate images using models trained before and after the exclusion of the top-1000 positive influencers identified by DAS and D-TRAK, using the same random seeds on ArtBench-2. We utilize 100 timesteps and a projection dimension of $k=32768$ to identify the top-1000 influencers. Additionally, we conduct a baseline experiment where 1000 training images are randomly removed before retraining. The results of this experiment are visualized in Figure 1.
Figure 1: We conduct an visualization experiment to explore DAS effectiveness. We retrain the model after deleting 1000 most influential training samples detected by D-TRAK and DAS. We also add an baseline that randomly delete 1000 samples. From Left to Right, they are the original generated image and generated by retrained model with random deletion, D-TRAK deletion and DAS deletion.
@@ 5 Conclusion
In this paper, we introduce the Diffusion Attribution Score (DAS) to address the existing gap in data attribution methodologies for generative models. We conducted a comprehensive theoretical analysis to elucidate the inherent challenges in applying TRAK to diffusion models. Subsequently, we derived DAS theoretically based on the properties of diffusion models for attributing data throughout the entire generation process. We also discuss strategies to accelerate computations to extend DAS to large-scale diffusion models. Our extensive experimental evaluations on datasets such as CIFAR, CelebA, and ArtBench demonstrate that DAS consistently surpasses existing baselines in terms of Linear Datamodeling Score (LDS) evaluation. This paper underscores the crucial role of data attribution in ensuring transparency and fairness in the use of diffusion models, especially when dealing with copyrighted or sensitive content. Looking forward, our future work aims to extend DAS to other generative models and real-world applications to further ascertain its effectiveness and applicability.
Appendix A Related Works
A.1 Data Attribution
The training data exerts a significant influence on the behavior of machine learning models. Data attribution aims to accurately assess the importance of each piece of training data in relation to the desired model outputs. However, methods of data attribution often face the challenge of balancing computational efficiency with accuracy. Sampling-based approaches, such as empirical influence functions (Feldman & Zhang, 2020), Shapley value estimators (Ghorbani & Zou, 2019; Jia et al., 2019), and datamodels (Ilyas et al., 2022), are able to precisely attribute influences to training data but typically necessitate the training of thousands of models to yield dependable results. On the other hand, methods like influence approximation (Koh & Liang, 2017; Schioppa et al., 2022) and gradient agreement scoring (Pruthi et al., 2020) provide computational benefits but may falter in terms of reliability in non-convex settings (Basu et al., 2021; Ilyas et al., 2022; Akyurek et al., 2022).
A.2 Data Attribution in Generative Models
The discussed methods address counterfactual questions within the context of discriminative models, focusing primarily on accuracy and model predictions. Extending these methodologies to generative models presents complexities due to the lack of clear labels or definitive ground truth. Research in this area includes efforts to compute influence within Generative Adversarial Networks (GANs) (Terashita et al., 2021) and Variational Autoencoders (VAEs) (Kong & Chaudhuri, 2021). Recently, Park et al. (2023) developed TRAK, a new attribution method that is both effective and computationally feasible for large-scale models. In the realm of diffusion models, earlier research has explored influence computation by employing ensembles that necessitate training multiple models on varied subsets of training data—a method less suited for traditionally trained models (Dai & Gifford, 2023). Journey TRAK (Georgiev et al., 2023) extends TRAK to diffusion models by attributing influence across individual denoising timesteps. Moreover, D-TRAK (Zheng et al., 2024) has revealed surprising results, indicating that theoretically dubious choices in the design of TRAK might enhance performance, highlighting the imperative for further exploration into data attribution within diffusion models. These studies are pivotal in advancing our understanding and fostering the development of instance-based interpretations in unsupervised learning contexts.
A.3 Application
Recent research has underscored the effectiveness of data attribution methods in a variety of applications. These include explaining model predictions (Koh & Liang, 2017; Ilyas et al., 2022), debugging model behaviors (Shah et al., 2023), assessing the contributions of training data (Ghorbani & Zou, 2019; Jia et al., 2019), identifying poisoned or mislabeled data (Lin et al., 2022), and managing data curation (Khanna et al., 2019; Liu et al., 2021; Jia et al., 2021). Additionally, the adoption of diffusion models in creative industries, as exemplified by Stable Diffusion and its variants, has grown significantly (Rombach et al., 2022; Zhang et al., 2023). This trend highlights the critical need for fair attribution methods that appropriately acknowledge and compensate artists whose works are utilized in training these models. Such methods are also crucial for addressing related legal and privacy concerns (Carlini et al., 2023; Somepalli et al., 2023).
Appendix B Proof of Equation 13
@@
In this section, we provide a detailed proof of Eq. 13. We utilize the Newton Method (Pregibon, 1981) to update the parameters in the diffusion model, where the parameter update $theta^{\prime}$ is defined as:
@@
Appendix C Proof of Equation 15
@@ Appendix D Algorithm of DAS
Appendix E Explanation about the choice of model output function in LDS evaluation
@@
Appendix F Implementation details
F.1 Datasets and Models
CIFAR10(32$\times$32). The CIFAR-10 dataset, introduced by Krizhevsky (2009), consists of 50,000 training images across various classes. For the Linear Datamodeling Score evluation, we utilize a subset of 1,000 images randomly selected from CIFAR-10’s test set. To manage computational demands effectively, we also create a smaller subset, CIFAR-2, which includes 5,000 training images and 1,000 validation images specifically drawn from the ”automobile” and ”horse” categories of CIFAR-10’s training and test sets, respectively.
@@ F.1 Datasets and Models
CelebA(64$\times$64). We selected 5,000 training samples and 1,000 validation samples from the original training and test sets of CelebA (Liu et al., 2015), respectively. Following the preprocessing method described by Song et al. (2021b), we first center-cropped the images to 140$\times$140 pixels and then resized them to 64$\times$64 pixels. For the CelebA experiments, we adapted the architecture to accommodate a 64$\times$64 resolution while employing a similar unconditional DDPM implementation as used for CIFAR-10. However, the U-Net architecture was expanded to 118.8 million parameters to better capture the increased complexity of the CelebA dataset. The hyperparameters, including the variance schedule, optimizer settings, and training protocol, were kept consistent with those used for the CIFAR-10 experiments.
@@
F.2 LDS evaluation setup
For the LDS evaluation, we construct 64 distinct subsets ${\mathbb{S}}_{m}$ from the training dataset ${\mathbb{S}}$, each constituting 50% of the total training set size. To ensure robustness and mitigate any effects of randomness, three separate models are trained on each subset using different random seeds. The LDS is computed by measuring the Spearman rank correlation between the model outputs and the attribution-based output predictions for selected samples. Specifically, we evaluate the Simple Loss $\mathcal{L}_{\text{Simple}}({\bm{z}},\theta)$ as defined in Eq 3 for samples of interest from both the validation and generation sets. To better approximate the expectation $\mathbb{E}_{t}$, we utilize 1,000 timesteps, evenly spaced within the range $[1,T]$. At each timestep, three instances of standard Gaussian noise ${\bm{\epsilon}}\sim\mathcal{N}(0,{\bm{I}})$ are introduced to approximate the expectation $\mathbb{E}_{\bm{\epsilon}}$. The calculated LDS values are then averaged across the selected samples from both the validation and generation sets to determine the overall LDS performance.
F.3 Baselines

### v4

@@ 5.3 Evaluation for Speed Up Techniques
The diffusion models used in our experiments are significantly complex, with parameter counts of 35.7M, 118.8M, and 25.5M respectively. These large dimensions pose considerable challenges in calculating the attribution score efficiently. To address this, we evaluate speed-up techniques as discussed in Section 4.3 on CIFAR-2. The results of these evaluations are reported in the Appendix E.4.
Normalization. We evaluate the normalization of gradients and residuals, as proposed in Eq. 17, to stabilize gradient variability across timesteps and enhance computational accuracy. By normalizing across generation before averaging, the performance for both DAS and D-TRAK improve (Table 4).
Number of timesteps. Computing DAS requires balancing effectiveness and computational efficiency, as more timesteps selected improve performance but increase costs. Experiment shows that while increasing timesteps enhances LDS results (Table 5), using 100 or 10 timesteps achieves comparable performance to 1000 timesteps with much lower computational demands.
Projection. We apply the projection technique to reduce gradient dimensions. As Johnson & Lindenstrauss (1984) discussed, higher projection dimensions better preserve inner products but increase computational costs. Figure 2 shows that LDS scores for both D-TRAK and DAS improve with increasing $k$ before plateauing. Based on these results, we set $k=32768$ as a default setting.
Compress Model Parameters. We explore techniques to reduce the gradient dimension by compressing. We conduct experiments by computing gradients on the up-block of U-Net, and the results in Table 6 show that this approach achieves competitive performance compared to the full model.
@@
Candidate Training Sample. Another technique to speed up the calculation involves reducing the number of training samples considered. We use CLIP to select the 1,000 training samples most similar to the target, compute attribution scores only for this candidate set (assigning 0 to others), and calculate the LDS, with results in Table 7 validating its effectiveness.
5.4 Evaluating Output Function Effectiveness
We define the output function as $f_{\text{DAS}}={\bm{\epsilon}}_{\theta}({\bm{x}}^{\text{gen}}_{t},t)$ to assess distribution shifts after data intervention. We argue that using Simple Loss introduces error as it involves the effect of data distribution during attribution. To validate this, we conduct a toy experiment using an unconditional DDPM trained on CIFAR-2 to generate 60 images. Each image is regenerated after removing 1,000 random training samples, producing 60 image pairs for which we compute $L^{2}$ distances. Denoising starts from timestep $T$, and we track average differences in loss and noise predictor outputs. Pearson correlation between $L^{2}$ distances and loss differences is 0.257, whereas for noise predictor differences, it reaches 0.485, indicating stronger alignment with image variations. Details are in Appendix H. This result aligns with intuition: if a model trained on cats and dogs generates a cat, removing all cat samples leads to a dog image under the same seed. Despite the drastic change, loss values may remain similar, as they reflect model optimization rather than generated content. Thus, loss differences fail to capture image changes, whereas noise predictor outputs effectively trace variations in diffusion model outputs.
5.5 Evaluating LDS for Various Attribution Methods
@@
In this section, we evaluate the performance of the Diffusion Attribution Score (DAS) against existing attribution baselines applicable to our experimental settings, as outlined by Zheng et al. (2024), with detailed explanations provided in Appendix E.3. Our primary focus is on comparing with post-hoc data attribution methods. To ensure fair comparisons, we limit the use of acceleration techniques for DAS to projection only, excluding others like normalization. The results on CIFAR, ArtBench and CelebA, presented in Tables 12, demonstrate that DAS consistently outperforms existing methods, where the result of CelebA is reported in the appendix in Appendix E.4 at Table 3.
@@
Compared to D-TRAK, DAS shows substantial improvements. On a validation set utilizing 100 timesteps, DAS achieves improvements of +9.33% on CIFAR-2, +8.39% on ArtBench-2, and +5.1% on CelebA. In the generation set, the gains continue with +5.01% on CIFAR-2, +5.78% on ArtBench-2, and +9.21% on CelebA. Notably, DAS also achieves significant improvements on larger datasets like CIFAR10 and ArtBench5, outperforming D-TRAK by +12.67% and +10.21% on validation sets and +7.58% and +8.92% on generation samples.
@@
Other methods generally underperform on larger datasets such as ArtBench5 and CIFAR10 compared to smaller datasets like CIFAR2 and ArtBench2. Conversely, our method performs better on ArtBench5 than on ArtBench2. Remarkably, our findings suggest that while with more timesteps for calculating gradients generally leads to a better approximation of the expectation $\mathbb{E}_{t}$, DAS, employing only a 10-timestep computation budget, still outperforms D-TRAK, which uses a 100-timestep budget in most cases, which underscores the effectiveness of our approach. Additionally, the modest improvements on CIFAR10 and CelebA may be attributed to the LDS setup for these datasets, which employs only one random seed per subset for training a model, whereas other datasets utilize three random seeds, potentially leading to inaccuracies in LDS evaluation. Another observation is that DAS performs better on the validation set than on the generation set. This could indicate that the quality of generated images may have a significant impact on data attribution performance. However, further investigation is needed to validate this hypothesis.
5.6 Counter Factual Visualization Evaluation
To more intuitively assess the faithfulness of DAS, we conduct a counter factual visualization experiment. We use different attribution methods, including TRAK, D-TRAK and DAS, to identify the the top-1000 positive influencers on 60 generated images on ArtBench2 and CIFAR2. We sample 100 timesteps and a projection dimension of $k=32768$ to identify the top-1000 influencers. These training samples identified by each method are subsequently removed and the model is retrained. Additionally, we conduct a baseline setting where 1,000 training images are randomly removed before retraining. We re-generate the images with same random seed and compare the $L^{2}$-Distance and CLIP cosine similarity between the origin and counter-factual images. For the pixel-wise $L^{2}$-Distance, D-TRAK yields values of 8.97 and 187.61 for CIFAR-2 and ArtBench-2, respectively, compared to TRAK’s values of 5.90 and 168.37, while DAS results in values of 10.58 and 203.76. In terms of CLIP similarity, DAS achieves median similarities of 0.83 and 0.71 for ArtBench-2 and CIFAR-2, respectively, which are notably lower than TRAK’s values of 0.94 and 0.84, as well as D-TRAK’s values of 0.88 and 0.77, demonstrating the effectiveness of our method. An illustration of the experiment is in Figure 1, where removing the influencers detected by DAS results in a biggest difference compared to baseline methods. A detailed result in box-plot is reported in Appendix G.
Figure 1: We conduct an visualization experiment to explore DAS effectiveness described in Sec 5.6. Removing the influential samples identified by DAS produces the most significant differences in the generated images after retraining the model. DAS is the most effective methods for attribution.
@@ 6 Conclusion
In this paper, we introduce the Diffusion Attribution Score (DAS) to address the existing gap in data attribution methodologies for generative models. We conducted a comprehensive theoretical analysis to elucidate the inherent challenges in applying TRAK to diffusion models. Subsequently, we derived DAS theoretically based on the properties of diffusion models for attributing data throughout the entire generation process. We also discuss strategies to accelerate computations to extend DAS to large-scale diffusion models. Our extensive experimental evaluations on datasets such as CIFAR, CelebA, and ArtBench demonstrate that DAS consistently surpasses existing baselines in terms of Linear Datamodeling Score evaluation. This paper underscores the crucial role of data attribution in ensuring transparency in the use of diffusion models, especially when dealing with copyrighted or sensitive content. Looking forward, our future work aims to extend DAS to other generative models and real-world applications to further ascertain its effectiveness and applicability.
7 Acknowledgment
This work was supported in part by the Start-up Grant (No. 9610680) of the City University of Hong Kong, Young Scientist Fund (No. 62406265) of NSFC, and the Australian Research Council under Projects DP240101848 and FT230100549.
Appendix A Proof of Equation 13
@@
In this section, we provide a detailed proof of Eq. 13. We utilize the Newton Method (Pregibon, 1981) to update the parameters in the diffusion model, where the parameter update $\theta^{\prime}$ is defined as:
@@
Appendix B Proof of Equation 15
@@ Appendix C Algorithm of DAS
Appendix D Explanation about the choice of model output function in LDS evaluation
@@
Appendix E Implementation details
(a) Val-10
(b) Val-100
(c) Gen-10
(d) Val-100
E.1 Datasets and Models
CIFAR10(32$\times$32). The CIFAR-10 dataset, introduced by Krizhevsky (2009), consists of 50,000 training images across various classes. For the Linear Datamodeling Score evaluation, we utilize a subset of 1,000 images randomly selected from CIFAR-10’s test set. To manage computational demands effectively, we also create a smaller subset, CIFAR-2, which includes 5,000 training images and 1,000 validation images specifically drawn from the ”automobile” and ”horse” categories of CIFAR-10’s training and test sets, respectively.
@@ E.1 Datasets and Models
CelebA(64$\times$64). We selected 5,000 training samples and 1,000 validation samples from the original training and test sets of CelebA (Liu et al., 2015), respectively. Following the preprocessing method described by Song et al. (2021b), we first center-cropped the images to 140$\times$140 pixels and then resized them to 64$\times$64 pixels. For the CelebA experiments, we adapted the architecture to accommodate a 64$\times$64 resolution while employing a similar unconditional DDPM implementation as used for CIFAR-10. However, the U-Net architecture was expanded to 118.8 million parameters to better capture the increased complexity of the CelebA dataset. The hyperparameter, including the variance schedule, optimizer settings, and training protocol, were kept consistent with those used for the CIFAR-10 experiments.
@@
E.2 LDS evaluation setup
Various methods are available for evaluating data attribution techniques, including the leave-one-out influence method (Koh & Liang, 2017; Basu et al., 2021) and Shapley values (Lundberg & Lee, 2017). In this paper, we use the Linear Datamodeling Score (LDS) (Ilyas et al., 2022) to assess data attribution methods. Given a model trained $\theta$ on dataset ${\mathbb{S}}$, LDS evaluates the effectiveness of a data attribution method $\tau$ by initially sampling a sub-dataset ${\mathbb{S}}^{\prime}\subset{\mathbb{S}}$ and retraining a model $\theta^{\prime}$ on ${\mathbb{S}}^{\prime}$. The attribution-based output prediction for an interested sample ${\bm{z}}^{\text{test}}$ is then calculated as:
where $\rho$ denotes the Spearman correlation and $\theta_{m}$ is the model trained on the $m$-th subset ${\mathbb{S}}^{m}$. For the LDS evaluation, we construct 64 distinct subsets ${\mathbb{S}}_{m}$ from the training dataset ${\mathbb{S}}$, each constituting 50% of the total training set size. For CIFAR-2, ArtBench-2, and ArtBench-5, three models per subset are trained using different random seeds for robustness, while for CIFAR-10 and CelebA, a single model is trained per subset. A validation set, comprising 1,000 samples each from the original test set and a generated dataset, serves as ${\bm{z}}^{\text{test}}$ for LDS calculations. Specifically, we evaluate the Simple Loss $\mathcal{L}_{\text{Simple}}({\bm{z}},\theta)$ as defined in Eq 3 for samples of interest from both the validation and generation sets. To better approximate the expectation $\mathbb{E}_{t}$, we utilize 1,000 timesteps, evenly spaced within the range $[1,T]$. At each timestep, three instances of standard Gaussian noise ${\bm{\epsilon}}\sim\mathcal{N}(0,{\bm{I}})$ are introduced to approximate the expectation $\mathbb{E}_{\bm{\epsilon}}$. The calculated LDS values are then averaged across the selected samples from both the validation and generation sets to determine the overall LDS performance.
(a) VAL-10
(b) VAL-100
(c) VAL-1000
(d) GEN-10
(e) GEN-100
(f) GEN-1000
E.3 Baselines